Due to erratic sales of its sole product-a high-capacity battery for laptop computers-PEM, Inc., has been experiencing financial difficulty for some time. The company's contribution format income statement for the most recent month is given below 

$ 402,000 Sales (13,400 units x $30 perr unit) Variable expenses Contribution margin Fixed expenses 201,000 201,000 223,500 $ (22,500) Net operating loss 

Required: 


1. Compute the company's CM ratio and its break-even point in unit sales and dollar sales. 

2. The president believes that a $6,300 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $89,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company's monthly net operating income? 

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $35,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? 

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by O.70 cents per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,400? 

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $55,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,000 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,000)?
Answer #1

1.CM ratio = CM/Sales

= 201,000/402,000

=50%

Break point in sales = Fixed costs/CM per unit

= 223,500/15

= 14,900 units

Dollar Sales = 223,500/50%

=$ 447,000

2.Increase in income = 89,000*50% - 6,300

= $38,200

3.Net Income = (27-15)*26,800 – 223,500-35,000

= $63,100

4.Target Profit = $4,400

Fixed cost = 223,500

Desired CM = $227,900

CM per unit = (30-15-0.7) = $14.30

Units = 227,900/14.30

= 15,937.06 units

5a.New CM ratio = 18/30 = 60%

Break even point in Unit sales = (223,500+55,000)/18

= 15,472.22 units

Dollar Sales = 278,500/60%

= $464,166.67

b.Income statement

No Automation

%

Automation

%

Sales

600,000

100%

600,000

100%

Variable Expenses

300,000

50%

240,000

40%

Contribution Margin

300,000

50%

360,000

60%

Fixed costs

223,500

37.25%

278,500

46.42%

Net Income

76,500

12.75%

81,500

13.58%

c.Yes, automation should be done (higher income)